# Magic Boundaries of 3D Color Codes

### v3

@@ V Summary and discussion
In principle, it is possible to include additional types of nested boundaries, specifically, those with nested $S$-domain walls where the walls intersect at multiple points, e.g., forming a crossing. As detailed in Ref. [17], at these intersections, $e$-particles emerge and attach to the boundary. However, these boundaries can be considered derivatives of the nested boundaries we explore in this paper. Consequently, we have chosen not to include them in our count.
@@
Future directions along this line include the application of the magic boundaries to fault-tolerant non-Clifford logical gates, where the initial exploration has occurred in Ref. [13, 38] in the context of fractcal topological codes. Besides, a natural theoretical framework describing such magic boundaries which go beyond the Lagrangian subgroup formalism is the Lagrangian algebra, which has been used to describe gapped boundaries in non-Abelian topological orders [46] as well as the twisted smooth boundary in the 3D toric code [43] in the framework of higher category theory. While the Lagrangian subgroup is a subgroup of condensable Abelian excitations, the Lagrangian algebra correspond to summing over all condensable excitations in the Lagrangian subgroup. Since the $T$-domain wall $s^{(3)}_{1,2,3}$ and $S$-domain wall $s^{(2)}_{i,j}$ can be understood as condensation defects corresponding to condensation of $e$ particles (i.e., summing over the electric Wilson lines) according to Ref. [17], it is expected that the magic boundary can be described by the Lagrangian algebra and the framework of higher category theory similar to the study in Ref. [43].
@@ Appendix A Detailed description of the unfolding unitary transformation
This section provides an detailed explanation of the unfolding unitary transformation introduced in Section II.4 and Ref. [11].
@@
One can further shrink the purple cells. The stabilizers in Eq. (102) become the stabilizers shown in Fig. 2, and the model becomes a 3D toric code model on purple lattice, as introduced in Sec. II.4.
@@
The shaded area on the left represents the fine structure of an yellow edge after we apply the lattice deformation. There are four qubits correspond to one edge and each qubit are coupled with a $Z$-stabilizer, which are displyed as triangles. On the right hand side, we zoom into this area, and there are two more types of $Z$-stabilizers. Both are two body stabilizers and couple qubits on two different directions. To get rid of these redundant qubits, one could first change the basis of stabilizers by product some of them together to make a single qubit on this edge couple to all the four different plaquettes, and then disentangle and discard the other three qubits. The remaining model is a 3D toric code on the yellow lattice as introduced in Sec. II.4.
@@ B.2.1 The smooth boundary
This boundary is exatly the smooth ($m$) boundary we discussed in Sec II. By checking the commutation relation between the $\mathbb{Z}_{2}$ global symmetry operator and the boundary Hamiltonian terms, one can also conclude that the ungauged Ising-model boundary corresponding to the $m$-boundary preserves the $\mathbb{Z}_{2}$ global symmetry.
@@ B.2.2 The rough boundary
The 00-form symmetry breaking can also be undertood in the context of $\mathbb{Z}_{2}$ gauge theory. Condensation of $e$-particles (which corresponds to the symmetry charges of the 00-form symmetry) violates the charge conservation symmetry. Therefore, the corresponding $\mathbb{Z}_{2}$ global symmetry is spontaneously broken on this boundary.
@@ B.2.3 The folded boundary
One can check this boundary corresponds to the $(e_{1}e_{2},m_{1}m_{2})$-boundary discussed in Se. III.3.3.
@@ Appendix C A Minimal Model Example
We select the top and bottom boundaries to be the $Z$-boundaries and the remaining four boundaries to be the $X$-boundaries, as we defined in Sec. III.
@@
At the intersection between the $X$-boundary and the $Z$-boundary, no additional stabilizers are introduced beyond those discussed in Sec. III.1.
One can further verify that after conjugation of the transversal-$T$ gate, the $X$-boundaries map to the magic boundaries, which then cannot condense either the $X$-type or the $Z$-type excitations, as we discuss in Sec. IV.4.
@@
In conclusion, the 3D color code obtained is equivalent to three copies of the toric codes, featuring $(e_{1},e_{2},e_{3})$-boundaries at the top and bottom, and $(m_{1},m_{2},m_{3})$-boundaries on the four remaining sides, up to local unitaries and ancillary qubits. Analogous to Fig. 3(b), we have the following relation.
[…]

### v4

@@ 5 Summary and discussion
In principle, it is possible to include additional types of nested boundaries, specifically, those with nested $S$-domain walls where the walls intersect at multiple points, e.g., forming a crossing. As detailed in Ref. [17], at these intersections, $e$-particles emerge and attach to the boundary. However, these boundaries can be derived from the nested boundaries we explore in this paper. Consequently, we have chosen not to include them in our count.
@@
Future directions along this line include the application of the magic boundaries to fault-tolerant non-Clifford logical gates, where the initial exploration has occurred in Ref. [13, 35] in the context of fractcal topological codes. Besides, a natural theoretical framework describing such magic boundaries which go beyond the Lagrangian subgroup formalism is the Lagrangian (condensable) algebra, which has been used to describe gapped boundaries and domain walls in non-Abelian topological orders in two spatial dimensions [43, 44, 12, 45, 46]. In three and higher spatial dimensions, the Lagrangian (condensable) algebra framework can be generalized using higher category theory. [47] Recent studies have explored examples of boundaries and domain walls, including the twisted smooth boundary in the 3D toric code [40, 42, 41], as well as type-I, type-II, and type-III domain walls between 3D toric codes [48]. Since the $T$-domain wall $s^{(3)}_{1,2,3}$ and $S$-domain wall $s^{(2)}_{i,j}$ can be understood as condensation defects corresponding to condensation of $e$ particles (i.e., summing over the electric Wilson lines) according to Ref. [17], it is expected that the magic boundary can be described by the Lagrangian algebra and the framework of higher category theory similar to the study in Ref. [40].
@@ Appendix A Detailed description of the unfolding unitary transformation
This section provides an detailed explanation of the unfolding unitary transformation introduced in Section 2.4 and Ref. [11].
@@
One can further shrink the purple cells. The stabilizers in Eq. (101) become the stabilizers shown in Fig. 2, and the model becomes a 3D toric code model on purple lattice, as introduced in Section 2.4.
@@
The shaded area on the left represents the fine structure of an yellow edge after we apply the lattice deformation. There are four qubits correspond to one edge and each qubit are coupled with a $Z$-stabilizer, which are displayed as triangles. On the right hand side, we zoom into this area, and there are two more types of $Z$-stabilizers. Both are two body stabilizers and couple qubits on two different directions. To get rid of these redundant qubits, one could first change the basis of stabilizers by product some of them together to make a single qubit on this edge couple to all the four different plaquettes, and then disentangle and discard the other three qubits. The remaining model is a 3D toric code on the yellow lattice as introduced in Section 2.4.
@@ B.2.1 The smooth boundary
which are single qubit $X$ terms and two-body $Z$ terms that are entirely supported on the boundary qubits.
@@
This boundary is exatly the smooth ($m$) boundary we discussed in Sec 2. By checking the commutation relation between the $\mathbb{Z}_{2}$ global symmetry operator and the boundary Hamiltonian terms, one can also conclude that the ungauged Ising-model boundary corresponding to the $m$-boundary preserves the $\mathbb{Z}_{2}$ global symmetry.
@@ B.2.2 The rough boundary
The 0-form symmetry breaking can also be undertood in the context of $\mathbb{Z}_{2}$ gauge theory. Condensation of $e$-particles (which corresponds to the symmetry charges of the 0-form symmetry) violates the charge conservation symmetry. Therefore, the corresponding $\mathbb{Z}_{2}$ global symmetry is spontaneously broken on this boundary.
@@ B.2.3 The folded boundary
One can check this boundary corresponds to the $(e_{1}e_{2},m_{1}m_{2})$-boundary discussed in Se. 3.3.3.
@@ Appendix C A Minimal Model Example
We select the top and bottom boundaries to be the $Z$-boundaries and the remaining four boundaries to be the $X$-boundaries, as we defined in Section 3.
@@
At the intersection between the $X$-boundary and the $Z$-boundary, no additional stabilizers are introduced beyond those discussed in Section 3.1.
One can further verify that after conjugation of the transversal-$T$ gate, the $X$-boundaries map to the magic boundaries, which then cannot condense either the $X$-type or the $Z$-type excitations, as we discuss in Section 4.5.
@@
Figure 13: Unfolding the 3D color code to three copies of the 3D toric codes, featuring $(e_{1},e_{2},e_{3})$-boundaries at the top and the bottom, and $(m_{1},m_{2},m_{3})$-boundaries on the four remaining sides.
@@
In conclusion, the 3D color code obtained is equivalent to three copies of the toric codes, featuring $(e_{1},e_{2},e_{3})$-boundaries at the top and bottom, and $(m_{1},m_{2},m_{3})$-boundaries on the four remaining sides, up to local unitaries and ancillary qubits. Analogous to Fig. 3(b), we show the relation in Fig. 13.